# Ultrafast Sampling-based Kinodynamic Planning via Differential Flatness

### 17 Mar 2026 v1

@@ V Technical Approach
In this section, we present our approach to kinodynamic planning by showing that the flat output evolves as a linear system (Sec. V-A), and hence, allows us to convert Problem 1 from the original state space $\cal X$} to a flat state space (Problem 2 in Sec. V-B). Problem 2 is, in turn, solved by our sampling-based kinodynamic planning with our FlatExtend subroutine in Sec. V-C, where closed-form solutions of the BVP and dynamics integration problems enables efficient parallelized forward kinematics and collision checking (Sec. V-D). Finally, we discuss trajectory simplification and optimality of our approach in Secs. V-E and V-F, respectively.
@@ V-A Flat Output Dynamics as a Linear System
Consider the $n$-dimensional flat output $\mathbf{y}$ in Def. 1, that can be used to recover the state $\mathbf{x}$ and control input $\mathbf{u}$ via Eq. (7). Let $\mathbf{w}\in\mathbb{R}^{n}$ be a pseudo-control input, defined as the $r$th derivative of the output $\mathbf{y}$:
$\mathbf{w}=\mathbf{y}^{(r)},\quad r=\max(l,m),$ (14)
where the derivative orders $l,m$ are defined in (7). Let $\mathbf{z}~\in~{\cal Z}~=~\mathbb{R}^{rn}$ be the flat state, defined as:
@@ V-B Sampling-based Kinodynamic Planning in Flat State Space
To solve Problem 2, we propose a sampling-based kinodynamic planning technique in the flat state space $\cal Z$} in Alg. 1, which is consistent with most existing sampling-based motion planners [karaman2010kinorrt, webb2013kinodynamicRRT, li2016sst]. Leveraging the linear dynamics (16), we develop an efficient FlatExtend subroutine (see Sec. V-C) to construct a graph or tree $\mathbb{G}_{\mathbf{z}}=(\mathbb{V}_{\mathbf{z}},\mathbb{E}_{\mathbf{z}})$ in the flat state space $\cal Z$}, where $\mathbb{V}_{\mathbf{z}}$ and $\mathbb{E}_{\mathbf{z}}$ denote the sets of nodes and edges, respectively, by solving the BVP and dynamics propagation problems for a polynomial local “flat” path $\mathbf{z}_{loc}(t)$. The resulting trajectory $\boldsymbol{\sigma}_{\mathbf{z}}(t)$ is a continuous piecewise-polynominal, consisting of $M$ local “flat” paths found on $\mathbb{G}_{\mathbf{z}}$:
@@
which is possible because the pseudo-control input $\mathbf{w}$ is the $r$-th order derivative of $\mathbf{y}$ with $r=\max(l,m)$ in Eq. (14).
Input: Initial state $\mathbf{x}_{s}$, initial control $\mathbf{u}_{s}$, goal region $\cal G$} with control input $\mathbf{u}_{\cal G}$, maximum iterations NN
@@ 4 $\mathbf{z}_{0}\leftarrow(\mathbf{y}_{0},\dot{\mathbf{y}}_{0},\ldots,\mathbf{y}^{(r-1)}_{0})$
10 $\mathbb{G}_{\mathbf{z}}\leftarrow$ FlatExtend($\mathbb{G}_{\mathbf{z}}$)
 /* Return $\boldsymbol{\sigma}(t)$ if reaching $\mathbb{G}_{\mathbf{z}}$ */
@@ 14    $\mathbf{u}(t)\leftarrow\boldsymbol{\beta}(\boldsymbol{\sigma}_{\mathbf{z}}(t),\mathbf{w}(t))$
Algorithm 1 Sampling-based Kinodynamic Motion Planning via Differential Flatness
V-C FlatExtend Subroutine on Flat Output Space
The goal of the FlatExtend subroutine, described in Alg. 2, is to find a collision-free dynamically feasible local “flat” path $\mathbf{z}_{loc}(t),t\in[0,T]$ with a time duration $T$, that connects an existing node $\mathbf{z}_{0}\in\mathbb{V}$ to a new node $\mathbf{z}_{f}$. A motion $\mathbf{z}_{loc}(t)$ can be generated by either sampling $\mathbf{z}_{f}$ and solving a BVP problem (Sec. V-C1) or sampling a pseudo-control input $\mathbf{w}_{0}$ and integrating the flat state dynamics (Sec. V-C2). For either case, we show that a closed-form expression of $\mathbf{z}_{loc}(t)$ can be obtained for parallelized collision checking in VectorizedCC subroutine in Sec. V-D. If $\mathbf{z}_{loc}(t)$ is valid, the node $\mathbf{z}_{f}$ is added to $\mathbb{V}_{\mathbf{z}}$ while the edge $(\mathbf{x}_{0},\mathbf{x}_{f})$ associated with the local “flat” path $\mathbf{z}_{loc}(t)$ is added to $\mathbb{E}_{\mathbf{z}}$.
@@ 3   Pick an existing node $\mathbf{Z}_{0}\in\mathbb{V}_{\mathbf{z}}$
4 $\mathbf{z}_{loc}(t)\leftarrow$ Eq. (24) with a sampled $T$ or an optimal $T$ from (25).
@@ V-C1 Solving the BVP Problem in Closed Forms
where $\mathbf{R}\succ 0$ is a user-defined positive definite weight matrix, and $\rho$ controls the trade-off between the control effort and the time it takes to finish the trajectory. We formulate the following Linear Quadratic Minimum Time (LQMT) problem [Verriest1991QuadMinTime] to solve for our local “flat” path $\mathbf{z}_{loc}(t)$:
@@
where the matrices $\mathbf{A}$ and $\mathbf{B}$ are defined in (16). Let us define $\mathbf{d}_{T}=\mathbf{z}_{f}-e^{\mathbf{A}T}\mathbf{z}_{0}$ and the Grammian matrix $\mathbf{G}_{T}=\int_{0}^{T}e^{\mathbf{A}t}\mathbf{B}\mathbf{R}^{-1}\mathbf{B}^{T}e^{\mathbf{A}^{\top}t}dt$}. The time duration $T$ can be either a) set to a fixed or sampled value or b) optimized for a minimum-time trajectory, as shown next.
@@
If we have the freedom to choose the duration $T$, we can solve the following equation for a minimum time [Verriest1991QuadMinTime]:
@@
For an arbitrary value of the pseudo-control order $r$, the minimum-time condition (25) can be solved for a positive $T$ using a numerical root-finding solver [nocedal1999numerical], that is amenable to simd parallelization such as L-BFGS [zhu1997lbfgsb]. More notably, many common systems such as manipulators (Example 1) or unicycles (Example 2) have $r=2$ while systems with higher $r$, such as quadrotors (Example 3), can reduce the pseudo-control order to $r=2$ to simplify motion planning in practice, as shown in [liu2017search]. For such cases, Example 4 below shows the optimal local “flat” path, pseudo-control input, and optimal duration, where the condition (25) is equivalent to solving a $4$th-order polynomial in (27) for a positive root, which can also be done in closed forms. As a result, solving for an optimal time $T$ is suitable for simd parallelization.
@@
The detailed derivation can be found in Sec. VIII-A.
@@ V-C2 Dynamics Propagation in Closed-Forms
Instead of sampling states and solving a BVP, another common approach for extending the graph/tree $\mathbb{G}_{\mathbf{z}}$ is to sample the pseudo-control input $\mathbf{w}_{loc}(t)=\mathbf{w}_{0}$ and the duration $T$, leading to a closed-form local path:
@@
where $\mathbf{y}_{loc}(t)=\frac{\mathbf{w}_{0}}{r!}t^{r}+\sum_{i=0}^{r-1}\frac{\mathbf{y}^{(i)}_{0}}{i!}t^{i}$. These local “flat” paths can be used to in tree-based kinodynamic planners such as Stable Sparse RRT [li2016sst] and are also suitable for parallelized forward kinematics and collision checking in Sec. V-D.
@@ V-D Vectorized Collision Checking
for $i=0,\ldots,\left\lceil\frac{N}{K}\right\rceil$. For each batch $\mathbf{b}_{i}$, we obtain a set of samples $\{\mathbf{x}_{loc}(t_{j})\}_{t_{j}\in\mathbf{b}_{i}}$ and perform fast parallelized collision checking via simd instructions, as illustrated in Fig. 2(b). The batch size $K$ is the number of floats that a simd register can store, e.g., $K=8$ for the commonly used AVX2 instruction set. The robot’s geometry is represented by a set of spheres, whose poses are calculated via forward kinematics for multiple states, and checked against the obstacles for collisions using simd parallelism. If any state in batch $\mathbf{b}_{i}$ leads to collision with an obstacle, we terminate the collision checking subroutine early and move on to other local paths. Similarly, other constraints such as state and control limits can be checked in parallel for each batch $\mathbf{b}_{i}$ to validate $\mathbf{z}_{loc}(t)$.
V-E Trajectory Simplification
Our kinodynamic planning approach in Sec. V-B returns a piecewise-polynomial trajectory: $\boldsymbol{\sigma}_{\mathbf{z}}(t)=\left\{(\mathbf{z}_{i}(t),t_{i})\right\}^{M}_{i=1}$, for $0=t_{0}<t_{1}<\ldots<t_{M}$ where the polynomial $\mathbf{z}_{i}(t),t\in[t_{i-1},t_{i}]$ corresponds to the pseudo-control $\mathbf{w}_{i}(t)$ and time duration $T_{i}=t_{i}-t_{i-1}$. Due to the sampling-based nature of our approach, the trajectory might contain unnecessary local paths and often requires further simplification, e.g., by looking for shortcuts between the nodes. For completeness, a simple simplification scheme is provided in Alg. 4 such that if there exists a collision-free trajectory $\mathbf{z}_{ij}(t)$ that connects two nodes $\mathbf{z}_{i}(t_{i-1})$ and $\mathbf{z}_{j}(t_{j})$, the local paths $\mathbf{z}_{k}$, $i\leq k\leq j$ in between can simply be bypassed by $\mathbf{z}_{ij}(t)$. We note that our approach is general and therefore, compatible with other complex simplification methods for sampling-based motion planning, such as [geraerts2007creating, hauser2010trajsim].
V-F Optimality
Let ${\cal C}^{*}=\min_{\boldsymbol{\sigma}_{\mathbf{z}}(t)}{\cal C}(\boldsymbol{\sigma}_{\mathbf{z}}(t),\mathbf{w}(t),T)$ be the minimum cost, defined in Eq. (20), over all collision-free trajectories that solve Problem 2. A trajectory $\boldsymbol{\sigma}_{\mathbf{z}}^{*}(t)$ is called optimal if it achieves the minimum cost $\cal C$$^{*}$. As we focus on solving for a closed-form local path and enabling fast trajectory validation via parallelization, our FlatExtend subroutine is compatible with common sampling-based kinodynamic motion planners such as [webb2013kinodynamicRRT, li2016sst]. Therefore, the optimality guarantees of the resulting trajectory $\boldsymbol{\sigma}_{\mathbf{z}}(t)$ on the flat state space $\cal Z$} depend on the optimality guarantees of the specific planner that integrates our FlatExtend subroutine. Furthermore, how optimality guarantees on $\cal Z$} translate to optimality guarantees in the original state space $\cal X$} is an interesting research question that we would like to leave for future work.
@@ V-F Optimality
    /* Bypass unnecessary motions */

### 21 Apr 2026 v2

@@ V Technical Approach
In this section, we present our kinodynamic planning framework, FLASK, by showing that the flat output evolves as a linear system (Sec. V-A), and hence, allows us to convert Problem 1 from the original state space $\cal X$} to a flat state space (Problem 2 in Sec. V-B). In the flat state space, we develop primitive subroutines, FlaskExtend for planning graph construction in Sec. V-C and FlaskCC for parallelized forward kinematics and collision checking in Sec. V-D, that work with any sampling-based planning method. These subroutines are the core of our framework in Alg. 1, which effectively turns any geometric motion planner into a kinodynamic version. As a result, instead of being tied to a specific planner, our general framework gives rise to a new class of ultrafast kinodynamic planners for a broad class of differentially flat robot systems. Finally, we discuss trajectory postprocessing in our approach in Sec. V-E.
@@ V-A Flat Output Dynamics as a Linear System
Consider the $n$-dimensional flat output $\mathbf{y}$ in Def. 1, that can be used to recover the state $\mathbf{x}$ and control input $\mathbf{u}$ via Eq. (7). Let $\mathbf{w}\in{{\cal W}\subset\mathbb{R}^{n}}$ be a pseudo-control input, defined as the $r$th derivative of the output $\mathbf{y}$:
$\mathbf{w}=\mathbf{y}^{(r)},\quad r{{\;\geq\;}}\max(l,m),$ (14)
where the derivative orders $l,m$ are defined in Eq. 7. Let $\mathbf{z}~\in~{\cal Z}\subset\mathbb{R}^{rn}$ be the flat state, defined as:
@@ V-B Sampling-based Kinodynamic Planning in Flat State Space
To solve Problem 2, we develop primitive subroutines in the flat state space $\cal Z$} for a sampling-based kinodynamic planning framework proposed in Alg. 1, which is consistent with most existing sampling-based motion planners [karaman2010kinorrt, webb2013kinodynamicRRT, li2016sst]. Leveraging the linear dynamics (16), we develop an efficient FlaskExtend subroutine (see Sec. V-C) to construct a graph or tree $\mathbb{G}_{\mathbf{z}}=(\mathbb{V}_{\mathbf{z}},\mathbb{E}_{\mathbf{z}})$ in the flat state space $\cal Z$}, where $\mathbb{V}_{\mathbf{z}}$ and $\mathbb{E}_{\mathbf{z}}$ denote the sets of nodes and edges, respectively, by solving the BVP and dynamics propagation problems for a polynomial local “flat” path $\mathbf{z}_{loc}(t)$. The resulting trajectory $\boldsymbol{\sigma}_{\mathbf{z}}(t)$ is a continuous piecewise-polynominal, consisting of $M$ local “flat” paths found on $\mathbb{G}_{\mathbf{z}}$:
@@
which is possible because the pseudo-control input $\mathbf{w}$ is the $r$-th order derivative of $\mathbf{y}$ with $r{{\;\geq\;}}\max(l,m)$ in Eq. (14).
Input: Initial state $\mathbf{x}_{s}$, initial control $\mathbf{u}_{s}$, goal region $\cal G$} with control input $\mathbf{u}_{\cal G}$, maximum iterations II
@@ 4 $\mathbf{z}_{0}\leftarrow(\mathbf{y}_{0},\dot{\mathbf{y}}_{0},\ldots,\mathbf{y}^{(r-1)}_{0})$
10 $\mathbb{G}_{\mathbf{z}}\leftarrow$ FlaskExtend($\mathbb{G}_{\mathbf{z}}$)
 /* Return $\boldsymbol{\sigma}(t)$ if reaching ${\cal G}_{\mathbf{z}}$ */
@@ 14    $\mathbf{u}(t)\leftarrow\boldsymbol{\beta}(\boldsymbol{\sigma}_{\mathbf{z}}(t),\mathbf{w}(t))$
Algorithm 1 FLASK: Sampling-based Kinodynamic Motion Planning via Differential Flatness
V-C FlaskExtend Subroutine on Flat Output Space
The goal of the FlaskExtend subroutine, described in Alg. 2, is to find a collision-free dynamically feasible local “flat” path $\mathbf{z}_{loc}(t),t\in[0,T]$ with a time duration $T$, that connects an existing node $\mathbf{z}_{0}\in\mathbb{V}$ to a new node $\mathbf{z}_{f}$. A motion $\mathbf{z}_{loc}(t)$ can be generated by either sampling $\mathbf{z}_{f}$ and solving a BVP problem (Sec. V-C1) or sampling a pseudo-control input $\mathbf{w}_{0}$ and integrating the flat state dynamics (Sec. V-C2). For either case, we show that a closed-form expression of $\mathbf{z}_{loc}(t)$ can be obtained for parallelized collision checking in FlaskCC subroutine in Sec. V-D. If $\mathbf{z}_{loc}(t)$ is valid, the node $\mathbf{z}_{f}$ is added to $\mathbb{V}_{\mathbf{z}}$ while the edge $(\mathbf{x}_{0},\mathbf{x}_{f})$ associated with the local “flat” path $\mathbf{z}_{loc}(t)$ is added to $\mathbb{E}_{\mathbf{z}}$.
@@ 3   Pick an existing node $\mathbf{Z}_{0}\in\mathbb{V}_{\mathbf{z}}$
4 $\mathbf{z}_{loc}(t)\leftarrow$ Eq. (25) with a sampled $T$ or an optimal $T=T^{*}$ from (26).
@@ V-C1 Solving the BVP Problem in Closed Forms
where $\mathbf{R}\succ 0$ is a user-defined positive definite weight matrix, and $\rho$ controls the trade-off between the control effort and the time it takes to finish the trajectory. We formulate the following Linear Quadratic Minimum Time (LQMT) problem [Verriest1991QuadMinTime] to solve for our local “flat” path $\mathbf{z}_{loc}(t)$ and control $\mathbf{w}_{loc}(t)$:
@@
where the matrices $\mathbf{A}$ and $\mathbf{B}$ are defined in (16). Let us define $\mathbf{d}_{T}=\mathbf{z}_{f}-e^{\mathbf{A}T}\mathbf{z}_{0}$ and the Gramian matrix
$\mathbf{G}_{T}=\int_{0}^{T}e^{\mathbf{A}t}\mathbf{B}\mathbf{R}^{-1}\mathbf{B}^{T}e^{\mathbf{A}^{\top}t}dt.$ (22)
The duration $T$ can be either a) set to a fixed or sampled value or b) optimized for a minimum-time trajectory, as shown next. Given a sample $\mathbf{z}_{f}$, the cost ${\cal C}_{loc}$ of the local path $\mathbf{z}_{loc}(t)$ is commonly used to choose an existing node $\mathbf{z}_{0}$ on the graph $\mathbb{G}_{\mathbf{z}}$, e.g., those in the neighborhood of $\mathbf{z}_{f}$ with ${\cal C}_{loc}$ smaller than a threshold $\zeta$ (see Sec. VI for how to choose the value of $\zeta$ as the number of nodes $N=|\mathbb{V}_{\mathbf{z}}|$ increases).
@@
If we have the freedom to choose the duration $T$, we can solve the following equation for a minimum time $T=T^{*}$ [Verriest1991QuadMinTime]:
@@
For an arbitrary value of the pseudo-control order $r$, the minimum-time condition (26) can be solved for a positive $T^{*}$ using a numerical root-finding solver [nocedal1999numerical], that is amenable to simd parallelization such as L-BFGS [zhu1997lbfgsb]. More notably, many common systems such as manipulators (Example 1) or unicycles (Example 2) have $r=2$ while systems with higher $r$, such as quadrotors (Example 3), can reduce the pseudo-control order to $r=2$ to simplify motion planning in practice, as shown in [liu2017search]. For such cases, Example 4 below shows the optimal local “flat” path, pseudo-control input, and optimal duration, where the condition (26) is equivalent to solving a $4$th-order polynomial in (28) for a positive root, which can also be done in closed forms. As a result, solving for an optimal time $T=T^{*}$ is suitable for simd parallelization.
@@
The detailed derivation can be found in Sec. X-A.
Remarks
We note that the closed-form polynomial solution (25) of the LQMT problem (21) is derived without considering additional constraints such as collision avoidance, velocity and acceleration limits. In our approach, such constraints are enforced via collision checking in Sec. V-D. If the trajectory $\mathbf{z}_{loc}(t)$ in (25) from $\mathbf{z}_{0}$ to $\mathbf{z}_{f}$ violates a constraint, it will be considered invalid. However, it does not necessarily mean that $\mathbf{z}_{f}$ is unreachable from $\mathbf{z}_{0}$. If there exists a constrained optimal trajectory $\mathbf{z}^{*}_{loc}(t)$, our Theorem 1 in Sec. VI shows that as the number of samples $N=|\mathbb{V}_{\mathbf{z}}|$ goes to infinity, the probability of having a piecewise-polynomial trajectory that is close to $\mathbf{z}^{*}_{loc}(t)$ (see Def. 2) tends to $1$, i.e., $\mathbf{z}_{f}$ can still be reached via a sequence of nodes with high probability.
Interestingly, recent results from optimal control [beaver2024optimal] find an optimal piecewise-polynomial local path under constraints for differentially flat systems by carefully switching between modes, each of which corresponds to a polynomial motion primitive. This approach solves an optimality condition on the cost function and the constraints to find an optimal mode schedule and is shown to generate an optimal trajectory for low-DOF systems with simple closed-form constraints in a few milliseconds. However, it is challenging to design such a mode-switching mechanism for high-DOF robots with complex geometry due to complicated forward kinematics, especially when the local path generation time is often restricted to the nanosecond range. Intriguingly, from a sampling-based perspective, our Theorem 1 (see Sec. VI) shows that such an optimal piecewise-polynomial local path can be found on our planning graph with high probability, i.e., a sequence of nodes on our graph is equivalent to the mode-switching schedule from [beaver2024optimal]. Nevertheless, this is an exciting direction for generating a constrained local path $\mathbf{z}_{loc}(t)$ in our framework that we leave for future work.
@@ V-C2 Dynamics Propagation in Closed-Forms
As BVP problems are challenging to solve, most existing work [li2016sst] resorts to dynamics propagation, where the planning tree is extended by sampling a constant control input and integrating the robot dynamics over small time steps using numerical methods. While our BVP closed-form solution in Sec. V-C1 is the main focus of our paper, this section shows that transforming the kinodynamic planning problem to the flat state space (Problem 2) also leads to closed-form dynamics propagation, and hence gets rid of numerical approximations with potentially high accumulated errors.
Instead of sampling the original control input, we sample the pseudo-control input $\mathbf{w}_{loc}(t)=\mathbf{w}_{0}$ and the duration $T$, leading to a closed-form local path:
@@
where $\mathbf{y}_{loc}(t)=\frac{\mathbf{w}_{0}}{r!}t^{r}+\sum_{i=0}^{r-1}\frac{\mathbf{y}^{(i)}_{0}}{i!}t^{i}$. These local “flat” paths can be used in tree-based kinodynamic planners such as Stable Sparse RRT [li2016sst] and are also suitable for parallelized forward kinematics and collision checking in Sec. V-D. Propagation-based planners tend to “wander” in the state space due to the suboptimality of random control with small time steps $T$. Interestingly, our BVP solutions can help by finding shortcuts to the goal rather than keep expanding the tree, and therefore reduce the planning time significantly as illustrated in Sec. VII-A.
@@ V-D Vectorized Collision Checking
for $i=0,\ldots,\left\lceil\frac{N}{K}\right\rceil$. For each batch $\mathbf{b}_{i}$, we obtain a set of samples $\{\mathbf{x}_{loc}(t_{j})\}_{t_{j}\in\mathbf{b}_{i}}$ and perform fast parallelized collision checking via simd instructions, as illustrated in Fig. 2(b). The batch size $K$ is the number of floats that a simd register can store, e.g., $K=8$ for the commonly used AVX2 instruction set. The robot’s geometry is represented by a set of SIMD-compatible primitive shapes such as spheres or capsules, whose poses are calculated via forward kinematics for multiple states and checked for collisions with obstacles using simd parallelism. If any state in batch $\mathbf{b}_{i}$ leads to collisions with an obstacle, we terminate the collision checking subroutine early and move on to other local paths.
Similarly, other constraints such as state and control limits can be checked in parallel for each batch $\mathbf{b}_{i}$ to validate $\mathbf{z}_{loc}(t)$. In general, the conversions (6) and (7) are nonlinear and might complicate the exact conversions of state and control limits on $(\mathbf{x},\mathbf{u})$ to corresponding flat state and pseudo-control limits on $(\mathbf{z},\mathbf{w})$. For many common robot systems, an upper bound on the conversions (7) can be obtained to estimate the limits in the flat state space as the flat output $\mathbf{y}$ is often part of the original state $\mathbf{x}$ (see Examples 12 and 3). For example, any limits on the joint angles and velocities of a manipulator in Example 1 can be directly translated to the flat state or any limits on the speed $v$ of a unicyle in Example 2 can be used to bound the first derivative of the flat out put $(\dot{x},\dot{y})$. Furthermore, since the flat output describes certain physical properties of the robot, e.g., the position and yaw angle of a quadrotor in Example 3, users can also directly specify the flat state and pseudo-control constraints, e.g., maximum linear velocity and yaw rates. For general nonlinear systems, this issue can be addressed by using large limits in the flat state space, and checking the local path $(\mathbf{x}_{loc}(t),\mathbf{u}_{loc}(t))$ in (30) against the original state and control limits on $(\mathbf{x},\mathbf{u})$ via sampling. However, there is a trade-off between the size of the flat state space and the risk of missing valid paths, which can be tuned based on the planner’s performance, e.g., the larger flat state space may lead to longer planning time but reduce the risk.
@@
    /* Bypass unnecessary motions if the trajectory $\mathbf{z}_{ij}(t)$ does not violates any constraints */